OpenStudy (zepp):

Determine the equation of the tangent to the graph of the function \(f(x)=x^3\) at the point \(\large \frac{1}{2}, \frac{1}{8}\)

OpenStudy (anonymous):

calculus?

OpenStudy (zepp):

Yeah

OpenStudy (anonymous):

let's suppose tangent is y=mx+b

OpenStudy (anonymous):

do u know how to get m?

OpenStudy (anonymous):

The tanget of a function is given by its derivate. The derivate of each point is the tangent of the original function on that given point. So taking the first derivative: \[f'(x) = 3x^2\] At the point x = 1/2 its value is \[f'(1/2) = { 3 \over 4}\]

OpenStudy (zepp):

So I differentiate \(\large y=x^3\)?

OpenStudy (anonymous):

yes just like what soati did

OpenStudy (anonymous):

according to the soati's statement m=3/4

OpenStudy (zepp):

Why did he substitute \(\frac{1}{2}\) into f'(x)?

OpenStudy (anonymous):

then you already have x and y for the point you want, just plug them in to find b

OpenStudy (anonymous):

because we want to find tangent at this point

OpenStudy (zepp):

Isn't f'(x) represent the slope of all points on the graph?

OpenStudy (zepp):

Ohh

OpenStudy (anonymous):

slope of the tangent is the derivative of the function at that point

OpenStudy (anonymous):

now what can we do for get b (y=3/4x+b)

OpenStudy (zepp):

Is the answer \(\large y=\frac{3}{4}x-\frac{1}{4}\)?

OpenStudy (anonymous):

yup

OpenStudy (zepp):

Thank you guys for clarifying this! :)

OpenStudy (zepp):

Also, this tangent crosses the main function at another place.

OpenStudy (zepp):

I need to find where; So I set up an equality; \(\large x^3=\frac{3}{4}x-\frac{1}{4}\)

OpenStudy (zepp):

I got \(x^3-3x+1=0\)

OpenStudy (zepp):

How do I factor that out?

OpenStudy (anonymous):

use ruffini. try dividing by x +1 Unless I got it wrong, the result should be 4x^2 - 4x + 1 So you will have (x+1)(4x^2 - 4x + 1) = 0

OpenStudy (anonymous):

\[x^3-\frac{3}{4}x+\frac{1}{4}=0 \\ x^3-\frac{1}{4}x-\frac{1}{2}x+\frac{1}{4}=0 \\ x(x^2-\frac{1}{4})-\frac{1}{2}(x-\frac{1}{2})=0 \\ x(x-\frac{1}{2})(x+\frac{1}{2})-\frac{1}{2}(x-\frac{1}{2})=0 \\ (x-\frac{1}{2})(x^2+\frac{1}{2}x-\frac{1}{2})=0\]

OpenStudy (anonymous):

by the way, your equation should be 4x^3 - 3x + 1 = 0

OpenStudy (zepp):

I see

OpenStudy (zepp):

Ok good! I got it, thank you so much! @mukushla I can't give out 2 medals :$, but thank you both! :)

OpenStudy (anonymous):

your welcome